One canned juice drink is 20% orange juice and another is 10% orange juice. How many liters of each should be mixed together in order to get 10L that is 11% orange juice
Answer by nerdybill(7384)

One canned juice drink is 20% orange juice and another is 10% orange juice. How many liters of each should be mixed together in order to get 10L that is 11% orange juice
.
Let x = liters of 20% orange juice
then
10-x = liters of 10% orange juice
.
.20x + .10(10-x) = .11(10)
.20x + 1 - .10x = 1.1
.20x - .10x = 0.1
.10x= 0.1
x = 0.1/.10
x = 1 liter of 20% orange juice
.
10% orange juice:
10-x = 10-1 = 9 liters